OpenStudy (anonymous):

(1 pt) What is the distance between the two points on a number line? Column A Column B 1. X = 4, Y = –17 2. P = 23, Q = 8 3. U = –1, V = –14 4. J = –3, K = 18 A. 13 B. 15 C. 21

OpenStudy (anonymous):

@HELPMEPLZ!!

OpenStudy (anonymous):

@lovergirl09876

OpenStudy (anonymous):

@Figureskater12

OpenStudy (anonymous):

HELP PLZ

OpenStudy (anonymous):

column a stops at j= k

OpenStudy (anonymous):

forget the k

OpenStudy (anonymous):

if u cant help its fine i understand

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

Let's take 6, 12 as our first example. We're trying to find out the difference between the two...take 12 minus 6 and you get a difference of 6 between them. Now the first problem, 4, -17...put the larger one first and subtract the smaller one. 4 minus negative 17 is the same as 4 plus 17 which is 21. 23-8 is 15, of course. -1-14 is -15, but the difference between something can't be negative, so it's just 15. 18-(-3) = 18+3 = 21. i found this online!

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

ok thx